Show the following operations on set(s)-

  • Complement of a set
  • Symmetric difference of two sets
1
Expert's answer
2021-11-15T17:55:52-0500

Let, U = {1,2,3......,10}

A = {1,3,5,7,9}

B = {2,3,4,5,6}

A' (the complement of the set A) = U - A = {2, 4, 6, 8, 10}

B' (the complement of the set B) = U - B = {1, 7, 8, 9, 10}

A â–³ B (the symmetric difference) = (A - B) U (B-A) = {1, 7, 9} U {2, 4, 6} = {1, 2, 4, 6, 7, 9}